1: Prepare a sales budget, including a schedule of expected cash collections.
2: prepare a merchandise purchase budget, including a schedule of expected cash disbursements for merchandise, and a selling and administrative budget.
3: prepare a cash budget.
MOST LIKELY NUMBERS AND ASSUMPTIONS SALES MANAGER PRIVATE INFORMATION October Most likely sales 600,000 910,000 475,000 385,000 PURCHASING MANAGER PRIVATE INFORMATION Most likely cost of merchandise as a % of sales Desired ending inventory as a percentage of next months cost of sales 45% 20% OPERATION MANAGER PRIVATE INFORMATION Most likely shipping expenses as a percent of sales Most likely other expenses as a percent of sales Salaries and wages (per month) Most likely advertising costs (per month) Most likely insurance costs (per month Depreciation expense (per month) 5.0% 8.0% 85,000 50,000 3,000 25,000 FINANCE MANAGER PRIVATE INFORMATION Percent sales collected in month of sale Percent sales collected in month after sale 30% 70% Percent of inventory purchases paid in month of purchase Percent of inventory purchases paid in month after purchase 50% 50% Percent of operating expenses paid in month of purchase Percent of operating expenses paid in month after purchase 100% 0% Desired minimum ending cash balance each month Borrow in increments of Monthly interest rate on borrowings (not 35,000 1,000 1% compounded) Other planned outlays of cash Capital expenditures 90 50,000 S
Sales Budget uly August September Quarter Sales Schedule of Cash Collections July August September Quarter June sales (A/R From Beginning Balance Sheet) July sales-Collected in Current Month July sales-Collected in Following Month August sales-Collected in Current Month August sales-Collected in Following Month September sales-Collected in Current Month Total Expected Cash Collections Total Sales for the Quarter A/R at the end of the Quarter(September Sales Still not Collected)
Inventory Purchase Budget August September October Sales Cost of Merchandise as % of Sales Budgeted Cost of Merchandise Sold Following Months COGS Desired Ending inventory % Desired Ending Inventory Dollars Budgeted Cost of Merchandise Sold Plus Desired Ending Inventory Total Inventory Needs Less Beginning Inventory Required Purchases Schedule of Expected Cash Disbursements Purchases July August September Quarter June Purchases (A/P From Balance Sheet) July Purchases August Purchases September Purchases Total Disbursements Cost of Merchandise Sold for the Quarter Ending Inventory at the end of the Quarter Ending AP at the end of the Quarter
Selling and Administrative Budget August September Quarter Sales Shipping as a Percentage of Sales Other Expenses as a Percentage of Sales Variable Expenses: Shipping Other Expenses Total Variable Expenses Fixed Expenses Salaries and Wages Advertising Prepaid Insurance Depreciation Total Fixed Expesnes Total Selling and Admin Expenses Less Noncash Items (Depreciation and Prepaid Ins) Total Cash Disbursements Prepaid Insurance That is Expensed During the Quarter Depreciation Expense Recognized During the Quarter
Cash Budget August Jul September Quarter Cash Balance: Beginning Add Cash Collections (From Sales Budget) Total cash Available Less Cash Disbursements For Inventory (From Purchasing Budget) For Operating Expenses (From Selling and Admin Budget) For Equipment For Cash Dividends For Interest (From Previous Months Borrowing) Total Cash Disbursements Excess (Deficiency)of Cash Financing Borrowing Repayment Total Financing Cash Balance: Ending Outstanding Loan Balance Interest on Borrowing (Due the Following Quarter)
Answer #1
Schedule of cash collections:
July August Sept Quarter
Collection:
June 0 0
July 180000 (30%) 420000 (70%) 600000
Aug 273000 (30%) 637000 (70%) 910000
Sept 142500 (30%) 142500
Total exp. Collections 180000 693000 779500 1652500
Total Sales for the Quarter = $1985000
A/R at the end of the Quarter (Sept sales still not collected) = 475000*0.7=$332500
Inventory Purchase budget:
July August Sept Quarter
Sales 600000 910000 475000
Cost of Merchandise as % of Sales 45% 45% 45%
Following Month's COGS 409500 213750 173250
Desired Ending Inventory % 20% 20% 20%
Desired Ending Inventory Dollars 81900 42750 34650
Budgeted Cost of Merchandise Sold 270000 409500 213750
Plus Desired Ending Inventory 81900 42750 34650
Total Inventory Needs 351900 452250 248400
Less Beginning Inventory 0 81900 42750
Required Purchases 351900 370350 205650 927900
Schedule of Expected Cash Disbursements - Purchases
July August Sept Quarter
For Purchases:
June 0 0
July 175950 175950 351900
Aug 185175 185175 370350
Sept 102825 102825
Total Disbursements 175950 361125 288000 825075
Cost of Merchandise Sold for the Quarter = 1985000 * 45% = $893250
Ending Inventory at the end of the Quarter = Ending of Sept = $34650
Ending A/P at the end of the Quarter = 50% of Sept purchases = 205650/2 = $102825
Schedule of Selling and Administrative Budget:
July August Sept Quarter
Sales 600000 910000 475000
Shipping as a Percentage of Sales 5% 5% 5%
Other Expenses as a Percentage of Sales 8% 8% 8%
Variable Expenses:
Shipping 30000 45500 23750 99250
Other Expenses 48000 72800 38000 158800
Total Variable Expenses 78000 118300 61750 258050
Fixed Expenses:
Salaries and Wages 85000 85000 85000 255000
Advertising 50000 50000 50000 150000
Prepaid Insurance 3000 3000 3000 9000
Depreciation 25000 25000 25000 75000
Total Fixed Expenses 163000 163000 163000 489000
Total Selling and Admin Expenses 241000 281300 224750 747050
Less: Non cash items (Dep. & Prepaid Ins.) 28000 28000 28000 84000
Total Cash Disbursements 213000 253300 196750 663050
Prepaid Insurance that is expensed during the quarter = 3000*3 = $9000
Depreciation Expense recognized during the Quarter = 25000 * 3 = $75000
CASH BUDGET :
July August Sept Quarter
Cash Balance Beginning 0 35000 35000 0
Add: Cash Collections 180000 693000 779500 1652500
Total Cash Available 180000 728000 814500 1652500
Less: Cash Disbursements:
For Inventory 175950 361125 288000 825075
For Operating Expenses 213000 253300 196750 663050
For Equipment 200000 90000 0 290000
For Cash Dividends 50000 0 0 50000
For Interest 0 411.63 421.49 833.12 (493950*1%*1/12)
Total Cash Disbursements 638950 704836.625 485171.49 1828958.1
Excess (Deficiency) of Cash -458950 23163.375 329328.51 -176458.12
Financing:
Borrowing 493950 11836.625 0 505786.63
Repayment 0 0 294328.51 294328.51
Total Financing 493950 11836.625 294328.51 211458.12
Cash Balance Ending 35000 35000 35000 35000
Outstanding loan balance = $505786.63 - 294328.51 = 211458.12
Outstanding interest on borrowing (due on the following quarter) = (505786.63*1%*1/12 = $421.49
(the repayment is being made at the end of sept after earning declaration)